OpenStudy (anonymous):

Express the required measurement in exact form. What is the degree measure of an arc 4 pi ft. long in a circle of radius 10 ft.?

OpenStudy (anonymous):

Arc is \[c=r \theta \] So you have \[4 \pi= 10 \theta \] So \[\theta=\frac{4 \pi}{10}=\frac{2 \pi}{5}\] In radians

OpenStudy (anonymous):

use theta=s/r so angle in radians is 4pi/10=2pi/5 Now angle in degrees is (180/pi)*(2pi/5)=360/5

OpenStudy (anonymous):

Meaning 72 degrees

OpenStudy (anonymous):

\[\theta=s/r\] \[\theta=2\pi/5\] \[\theta=(2\pi/5)(180/\pi)=72 degrees\]